OpenStudy (anonymous):

So would the derivative of y=Arcsin(3x/4) be [3/\sqrt{16-9x^2}\]?

OpenStudy (anonymous):

Just a quick note, you need to put a \ in front of the [ at the beginning for it to register as an equation. As for your question, your answer is correct. Good job!

OpenStudy (anonymous):

Thanks for the tip.

OpenStudy (anonymous):

Also, another handy notation to use is \frac{x}{y}, which registers as a fraction with whatever you type into x as the numerator and y as the denominator, so writing \frac{3}{\sqrt{16-9x^2}} will be seen as \[\frac{3}{\sqrt{16-9x^2}}\] This just makes some fractions easier to read than just using x/y for instance.

OpenStudy (anonymous):

ok great